you are in the parachute flying over the Jurassic island when you spot a dinosaur. if the parachute is 15,000 ft in the air and the distance from you to the dinosaur is 8,500ft what Angle of depression are you looking when you spotted the dinosaur
Found 3 solutions by Theo, ikleyn, math_tutor2020:
Answer by Theo(13342)

my diagram is shown below.



the vertical length if 1500
the horizontal length is 8500
the angle x is found by taking the arctangent of (8500/1500).
the angle is 29.54 degrees.

Answer by ikleyn(52945)

.
you are in the parachute flying over the Jurassic island when you spot a dinosaur.
if the parachute is 15,000 ft in the air and the distance from you to the dinosaur is 8,500ft,
what Angle of depression are you looking when you spotted the dinosaur?
~~~~~~~~~~~~~~~~~


The problem is posed incorrectly and is SELF-CONTRADICTORY.

The distance to the dinosaur MUST be greater than the height of the parachute.

In the problem, everything is in opposite.


Revise your problem and make it in correct way.



Answer by math_tutor2020(3820)


I agree with tutor Ikleyn. The phrasing "the distance from you to the dinosaur is 8,500ft" tells us that the hypotenuse is 8500 ft; but the hypotenuse must be the longest side of the right triangle.

If it was instead phrased as "the horizontal distance from you to the dinosaur is 8500 ft", then tutor Theo's drawing would apply.

However, Theo appears to have misinterpreted what an angle of depression is.
Angle x of his drawing is NOT the angle of depression.

The angle of depression results when you first look horizontally and then look down at your target.
Here's an example

Image Source
https://www.mathsisfun.com/definitions/angle-of-depression.html

We must subtract the x value that Theo got from 90 to get the angle of depression.
90 - x = 90 - 29.54 = 60.46
Or we could compute it like this
arctan(15000/8500) = 60.46 approximately